Elektra receives 75% of its electronic components from vendor A and the remaining 25% from vendor B. The percentage of defectives from vendors A and B are 190 and 2%, respectively. When a random sample of size 5 from a received lot is inspect ne defective unit is found. Determine the probability that the lot is received from vendor A. From vendor B. (Hint: The probability distribution of defective items in a sample is binomial.)
Answer #1

P(1 defective)=P(from A and 1 defective)+P(from B and 1 defective)

=0.75*5C1(0.01)1(0.99)4+0.25*5C1(0.02)1(0.98)4 =0.75*0.0480+0.25*0.0922 =0.0591

hence P(lot from vendor A given 1 defective)=P(from A and 1 defective)/P(1 defective)=0.75*0.0480/0.0591=0.6097

P(lot from vendor B given 1 defective)=P(from B and 1 defective)/P(1 defective)=0.25*0.0922/0.0591=0.3903